if there are 90 pizzas, and each person eats 1/6 of the pizzas--How many people are there?
Answer by Theo(13342)

each person eats 1/6 of a pizza.
6 people per pizza
90 pizzas
number of persons is equal to 6 * 90 = 540.

formula is:

number of people * amount of pizza eaten per person = number of pizzas eaten.

n = number of people
r = amount of pizza eaten per person
p = number of pizzas eaten

n * r = p

r = 1/6 because each person eats 1/6 of a pizza.

p = 90 (given)

n * r = p becomes:

n * (1/6) = 90

multiply both sides of this equation by 6 to get:

n = 90 * 6 which becomes:

n = 540

540 people eating 1/6 of a pizza each consume 90 pizzas (540 * 1/6 = 90)